Stock A: X = $24.50, s = 5.25 Stock B: X = $147.25, s = 12.25 Stock C: X $5.75, s 2.08 (a) Find the coefficient of variation for prices of these three stocks. (Round your answers to 2 decimal places.) Coefficient of Stock Stock A Stock B Stock C variation (b) Which stock has the greatest relative variation? Stock A Stock B Stock C
Answer #1

a) Stock A CV = (5.25/24.5)*100 = 21.43

Stock B CV = (12.25/147.25)*100 = 8.32

Stock C CV = (2.08/5.75)*100 = 36.17%

b) Stock C